OpenStudy (anonymous):

(wx - y)2 but it have to b put on a grid like polynomials

OpenStudy (espex):

Do you have to expand this expression?

OpenStudy (anonymous):

yes

OpenStudy (espex):

All you need to do is to FOIL the expression times itself: \[(wx-y)(wx-y)\]

OpenStudy (anonymous):

like I have to put them an a box n order

OpenStudy (espex):

I'm not sure I understand what you're needing to do, can you explain a little more or give me an example?

OpenStudy (anonymous):

sure (x + y)(x - y) answer is x squared y squared

OpenStudy (anonymous):

(0.1x + 0.4y)2

OpenStudy (espex):

Okay, so you will just do that exact same thing with this problem. If it helps, think of wx as a single 'x'.

OpenStudy (espex):

Again, the same process, you are squaring the whole (0.1x+0.4y), so you will have it multiplied by itself.

OpenStudy (anonymous):

is there a number I can reach yuh at

OpenStudy (espex):

No, but there is the thread we're working in. :)

OpenStudy (anonymous):

okay what about this (2x -1 over two )2

OpenStudy (espex):

If we are assuming that this is an exercise in expansion, you will do the same with this problem. \[(2x-\frac{1}{2})(2x-\frac{1}{2})\] Simply FOIL the terms to expand your factors.

OpenStudy (anonymous):

(x + 2)(x - 2)

OpenStudy (espex):

\((a+b)(a-b) = (a^2-b^2)\)